OpenStudy (anonymous):

Translations: if the home equation is f(x) = x^2, why does f(x) = (x-3)^2 move the equation to the right instead of the left?

OpenStudy (irishboy123):

\(y = x^2\) \(x = \pm \ \sqrt{ y}\) \(\bar x = x + 3\) \(\bar x = \pm \ \sqrt{ y} + 3\) \(\pm \ \sqrt{ y} = \bar x - 3 \) \(y = ( \bar x - 3)^2 \)

OpenStudy (irishboy123):

in general speak \(y = f(x)\) \(x = f^{-1}(y)\) \(\bar x = x + a\) \(\bar x = f^{-1}(y) + a\) \(f^{-1}(y) = \bar x - a\) \(y = f(\bar x - a)\)